# PyFi: Toward Pyramid-like Financial Image Understanding for VLMs via Adversarial Agents

### v1

@@ Abstract
This paper proposes PyFi, a novel framework for pyramid-like financial image understanding that enables vision language models (VLMs) to reason through question chains in a progressive, simple-to-complex manner. At the core of PyFi is PyFi-600K, a dataset comprising 600K financial question-answer pairs organized into a reasoning pyramid: questions at the base require only basic perception, while those toward the apex demand increasing levels of capability in financial visual understanding and expertise. This data is scalable because it is synthesized without human annotations, using PyFi-adv, a multi-agent adversarial mechanism under the Monte Carlo Tree Search (MCTS) paradigm, in which, for each image, a challenger agent competes with a solver agent by generating question chains that progressively probe deeper capability levels in financial visual reasoning. Leveraging this dataset, we present fine-grained, hierarchical, and comprehensive evaluations of advanced VLMs in the financial domain. Moreover, fine-tuning Qwen2.5-VL-3B and Qwen2.5-VL-7B on the pyramid-structured question chains enables these models to answer complex financial questions by decomposing them into sub-questions with gradually increasing reasoning demands, yielding average accuracy improvements of 19.52% and 8.06%, respectively, on the dataset. All resources of code, dataset and models are available at: https://github.com/AgenticFinLab/PyFi.
@@ 1 Introduction
Vision language models (VLMs) are capable of step-by-step reasoning, presented as Chain-of-Thought (CoT) CoT-neurips22, for image understanding in question answering (Q&A) across various scientific domains (MMStar-neurips24; MMMU-cvpr24; Visco-cvpr25), driven by the development of domain-specific benchmarks (VisNumBench-iccv25; Mvmath-cvpr25; MMMU-cvpr24; Codevision-arxiv25), step-wise annotations (MathShepherd-arxiv23; o1-coder-arxiv24; mutualreasoning-arxiv24), and domain-tailored VLMs (FinGPT-ijcai23; open-finllms-arxiv24; finlmm-r1-arxiv25). In contrast, advancements in the financial domain remain limited. This is because, in finance image understanding, the creation of CoT reasoning datasets, which are highly necessary for the evaluation and fine-tuning of VLMs, requires a high level of expertise and domain knowledge and is therefore scarce.
@@
To address these issues, previous literature (FinQA-emnlp21) collects Q&A samples from financial documents manually, a process that is prohibitively time-consuming and difficult to scale. As a result, subsequent works propose datasets by merging extensive existing financial datasets (FLUE-emnlp22; Finben-neurips24). These repurposed datasets may introduce accumuated biases and inconsistencies inherited from the original sources. More recently, benchmarks such as (Mme-finance-mm25; FCMR-acl25; FinMME-acl25; BizFinBench-arxiv25; CFBenchmark-MM-arxiv25; VisFinEval-arxiv25) leverage trained VLMs to generate Q&A pairs augmented with CoTs (FAMMA-arxiv25). However, these samples may contain factual errors or hallucinations, which can distort VLM evaluation and fine-tuning and thereby undermine their reliability in real-world financial applications, where precision and expert-level understanding are essential.
@@
PyFi-600K offers three key benefits for the interpretability and fine-tuning of VLMs in finance. First, the samples are organized in a pyramid structure, where the required image understanding capability increases progressively — from basic visual perception at the bottom level to complex financial decision-making at the top. This hierarchy, with capability levels ranging from one to six, enables comprehensive evaluation of VLMs across multiple aspects and supports ranking the capabilities required by VLMs in different financial themes. Second, samples across different levels are interconnected through question chains, demonstrating how questions at higher levels can be addressed step-by-step by leveraging information and reasoning from lower-level Q&A samples. Finally, it supports process supervision (PRM800K-iclr23), as each sample is assigned a reward score that reflects the correctness of its reasoning process in solving the financial problem.
@@ 2 Related Work
Financial Benchmark. The evaluation of Vision-Language Models (VLMs) in the finance domain has seen remarkable progress, from early text-based benchmarks (FinQA-emnlp21; FLUE-emnlp22; Finben-neurips24) to more recent efforts targeting financial image understanding (FCMR-acl25; FinTMMBench-mm25; FinChartBench-arxiv25; CFBenchmark-MM-arxiv25; VisFinEval-arxiv25; Fin-Fact-mm25; FinRAGBench-V-arxiv25; Mme-finance-mm25; FinMME-acl25; FAMMA-arxiv25; FinMR-iclr25). However, the question-answering datasets of most benchmarks, such as FinMME (Mme-finance-mm25), rely on human annotation, making them costly and difficult to scale. Besides, they fail to support comprehensive assessment of VLMs in the financial domain. However, we introduce an automatically synthesized dataset for the fine-grained, hierarchical evaluation of financial visual understanding across diverse image types and financial themes.
Financial Large Models. Inspired by the success of large models (LMs) in general domains, recent research has increasingly focused on developing specialized financial LMs to address tasks requiring deep domain expertise. Previous efforts primarily centered on text analysis using financial LMs (FinBERT-car23; Bloomberggpt-arxiv23; FinGPT-ijcai23; PIXIU-neurips23; Finr1-arxiv25). Recent financial VLMs (Finvis-gpt-ijcai23; FinTral-acl24; open-finllms-arxiv24; FinTab-LLaVA-kdd25) integrate visual inputs to facilitate step-by-step reasoning, thereby enabling broader financial decision-making. However, these models lack the expertise and interpretability needed for challenging financial themes, a limitation stemming from the absence of detailed, step-wise problem-solving processes grounded in finance-specific knowledge. In contrast, our dataset contains questions that illustrate how to address financial problems through a progression from simple to complex image understanding and from basic to increasingly deep financial knowledge.
Multi-Agent LM Framework. Recent advances in LMs have enabled Agentic AI to excel in real-world applications, with multiple LM-based agents now collaborating to perceive, learn, reason, and act, supporting intricate tasks such as simulation (Multi-agent-simulation-arxiv25; Multi-agent-simulator-acl25), reasoning (BeTop-neurips24; AgenticReasoning-neurips25; MultiAgentDebate-icml24), trading (Hedgeagents-www25; FinCon-neurips24; MountainLion-arxiv25), and data synthesis (systematic-simulation-arxiv24; STORM-BORN-acl25; AgentSGEN-arxiv25). However, to our knowledge, no existing method in finance leverages a multi-agent adversarial approach to synthesize data. In contrast, inspired by Generative Adversarial Networks (GANs) gan-neurips14, we present the first method that employs adversarial agents to automatically synthesize hierarchical, fine-grained financial datasets.
@@ 3 Framework
Figure 2: Overview of PyFi-adv: a challenger agent competes with a solver agent under the MCTS paradigm to generate question chains that progressively probe deeper capability levels in financial visual reasoning.
@@ 3.1 Financial Dataset in Pyramid Structure
Therefore, we observe the following three benefits for the finance domain:
@@
Step-wise Reward Score $r$ of each sample reflects its reliability; thus, the scores along a sample chain support process supervision (PRM800K-iclr23) in reasoning with financial image understanding. Using these scores, we enable the training of process reward models that act as verifiers for any financial VLM.
@@ 3.2 Automatic Synthesization with Adversarial Agents
Existing financial datasets rely solely on human annotation, while automated methods from other domains, such as rStar-Math (rStar-Math-arxiv25) for mathematical reasoning, fail to meet the requirements of finance, where interpretability and reliable reasoning are essential, along with increasing complexity, domain expertise, and hierarchical cognition. We achieve this by proposing a multi-agent adversarial mechanism, referred to as PyFi-adv, which constructs question chains, as illustrated in Figure 2
Specifically, PyFi-adv contains a challenger agent that is supported by an off-and-shelf VLM, denoted as $\psi$, and similarly, a solver agent $\phi$. Inspired by the confrontation in the game of Go used in AlphaGo (AlphaGo-nature), our two agents compete within a MCTS framework to automatically synthesize samples.
@@ 4.1 Setup
Table 1: Evaluation of 15 Pre-trained VLMs on PyFi-600K across $6$ capability levels, 55 complexity degrees. For simplicity, we only present the results from degrees 33, 44, and 55. The test set consists of $301$ samples, derived from an initial random selection of $1000$ instances followed by an additional filtering step. The full model names listed in the “Model” column are: InternVL3-38B, Claude-opus-4-1-20250805, DeepSeek-VL2, Hunyuan-Large-Vision, ERNIE-4.5-turbo-vl, Moonshot-V1-128k-Vision-Preview, Moonshot-V1-32k-Vision-Preview, Moonshot-V1-8k-Vision-Preview, Qwen3-VL-Plus, Qwen2.5-VL-72B-Instruct, Qwen2.5-VL-32B-Instruct, Qwen2.5-VL-7B-Instruct, and Qwen2.5-VL-3B-Instruct.
Model Overall PP DE CA PR LR DS 33 44 55 GPT-4.1 52.99 86.84 47.64 40.62 57.14 54.84 38.46 52.63 46.55 53.85 InternVL3 52.91 76.32 56.37 44.53 43.88 53.23 30.77 41.78 44.83 67.31 Claude-opus 64.70 80.92 62.74 66.02 58.16 72.58 32.69 66.12 55.60 73.08 GLM-4.5V 74.75 89.47 78.07 65.23 75.51 75.81 46.15 72.37 64.22 73.08 DS-VL2 45.18 76.32 44.58 36.33 42.35 42.74 19.23 34.54 42.24 46.15 Hunyuan-LV 59.72 90.79 58.96 36.72 63.27 69.35 51.92 50.00 57.76 80.77 ER-4.5-TVL 34.47 34.87 33.96 36.72 25.00 45.16 36.54 38.82 25.43 65.38 MV1-128k 54.57 80.92 61.08 39.06 41.84 58.87 38.46 46.71 43.97 57.69 MV1-32k 54.40 80.26 61.56 37.11 40.31 62.90 38.46 45.72 43.97 61.54 MV1-8k 54.90 80.26 61.79 39.84 42.35 58.06 38.46 48.03 43.97 55.77 Q3-VL-Plus 51.00 77.63 54.25 37.50 41.84 57.26 32.69 45.72 37.50 65.38 Q2.5-VL-72B 48.84 70.39 59.91 33.20 35.71 46.77 26.92 36.51 31.03 57.69 Q2.5-VL-32B 43.19 60.53 45.28 37.50 36.73 42.74 28.85 39.14 30.60 59.62 Q2.5-VL-7B 37.87 54.61 49.53 34.77 18.37 24.19 15.38 31.91 17.67 38.46 Q2.5-VL-3B 20.51 36.84 15.57 21.48 16.84 21.77 19.23 23.03 18.97 30.77
Datasets. PyFi-600K dataset consists of 62,660 sample chains with reward scores of 1. For evaluation, we proportionally sample 1,000 samples from each capability level to construct the test set across all cases. For supervised fine-tuning (SFT), we directly filter out samples for which more than 8 out of 15 VLMs produce the same answer even when the financial image is missing. This additional filtering step ensures the removal of samples that may have been memorized by VLMs during pre-training, a phenomenon referred to as data leakage.
@@ 4.3 Main Benefits and Insights
Hierarchical Evaluation of VLMs in Table 1 and Figure 3 enables a fine-grained analysis of how — and to what extent — VLMs meet the varying capability requirements of financial image understanding across different levels of complexity. The phenomenon of decreasing accuracy as the level increases in Table 1 indicates that existing VLMs can only meet the most basic image understanding in finance and are difficult to be used for complex financial decision-making. For instance, while models like GLM-4.5V perform well on basic perception questions (level 1), their accuracy drops considerably on decision support questions (level 6). This trend is consistent across most models, suggesting that greater efforts are needed to enhance VLMs’ capabilities in the finance domain. Thus, the results in Figure 3 demonstrate the importance of our question-chain fine-tuning: Qwen models learn to make accurate financial decisions by first addressing a series of sub-questions in a progression from simple to complex, a process aligned with human cognition.
@@
Tracing back from Decision Support to Lower-level Q&As, as illustrated in Figure 4, significantly improves the interpretability of VLMs in financial decision-making. In Figure 4, we first collect Level-6 samples that VLMs fail to answer correctly and analyze their question chains by evaluating VLMs on the questions at each lower level in the chain. In the finance domain, addressing a Level-6 financial decision-making question typically requires correctly answering sub-questions from lower levels. Thus, when an error occurs, as indicated by the red cross in the right sub-figure, PyFi-600K enables us to trace the problem-solving chain from DS to LR and eventually to PP, thereby diagnosing how an error in the DE leads to the final mistake. Such a tracing-back approach is crucial for understanding the limitations of VLMs in financial decision-making, as it reveals at which capability level a failure occurs. It is also essential for guiding how to adjust the reasoning chain to achieve accurate final decisions.
@@
Identifying the Most Frequent Failure Capability Levels, as presented in the left part of Figure 4, reveals that the key to improving VLMs’ performance in the financial domain lies in strengthening their capabilities at the CA (Calculation Analysis) level. Specifically, the pyramid-like structure and the question chain of PyFi-600K facilitates detailed error analysis, allowing us to identify specific levels where VLMs struggle. That is, most VLMs, even the advanced Claude-opus-4-1-20250805 and Qwen3-VL-Plus, fails to answer more than 40% of the question from CA, revealing that the calculation analysis is a critical bottleneck for them. More importantly, by comparing the error proportions between six capability levels, we can prioritize which capabilities to enhance first. For instance, since CA has the highest error proportion, focusing on improving calculation analysis skills in VLMs could lead to the most significant overall performance gains in financial decision-making tasks. In addition, this comparison also shows that, in the finance domain, analyzing numerical calculations from images is one of the most critical capabilities for accurate decision-making, even more important than logical reasoning, aligning with conclusions drawn by real-world finance experts  [^1] We consulted 15 experts from financial quantification, market simulation, investment practice, and academia to arrive at this conclusion..
@@
Addressing Financial Decision-Making with Level-wise CoT, as depicted in Figure 5, is not only reliable but also highly interpretable. Specifically, we compare the responses of Qwen2.5-VL-3B-Instruct and PyFi-QwenVL-3B — the latter fine-tuned on question chains— when addressing the financial problem. The Qwen2.5-VL-3B-Instruct produces a short and disorganized reasoning process, reflecting its limited financial reasoning capability. In contrast, PyFi-QwenVL-3B follows a clear, structured, and logical reasoning flow, which is essential for expert-level performance in finance. Mirroring the pyramid structure of PyFi-600K and the progressively increasing capability level of questions in the question chains, our model first addresses basic PP questions, then proceeds to DE and CA, and finally tackles high-level LR and DS. This level-wise CoT not only improves answer reliability but also provides a transparent and traceable reasoning path that users can easily follow and verify, making it especially valuable in financial contexts where stakes are high. Moreover, experts can inspect this professional decision-making process and provide detailed and direct feedback to further refine the model.
@@
Providing Process Supervision, in the form of a reward score for each sample in our PyFi-600K, enables research into training reliable verifier models, known as process reward models (PRMs), to facilitate checking the correctness of each step generated by financial VLMs. Even though, due to space limitations, this paper does not perform such training, we use the reward scores to filter out low-quality samples during both evaluation and fine-tuning, as discussed in the experimental setup.
5 Concluding Remarks

### v2

@@ Abstract
This paper proposes PyFi, a novel framework for pyramid-like financial image understanding that enables vision language models (VLMs) to reason through question chains in a progressive, simple-to-complex manner. At the core of PyFi is PyFi-600K, a dataset comprising 600K financial question-answer pairs organized into a reasoning pyramid: questions at the base require only basic perception, while those toward the apex demand increasing levels of capability in financial visual understanding and expertise. This data is scalable because it is synthesized without human annotations, using PyFi-adv, a multi-agent adversarial mechanism under the Monte Carlo Tree Search (MCTS) paradigm, in which, for each image, a challenger agent competes with a solver agent by generating question chains that progressively probe deeper capability levels in financial visual reasoning. Leveraging this dataset, we present fine-grained, hierarchical, and comprehensive evaluations of advanced VLMs in the financial domain. Moreover, fine-tuning Qwen2.5-VL-3B and Qwen2.5-VL-7B on the pyramid-structured question chains enables these models to answer complex financial questions by decomposing them into sub-questions with gradually increasing reasoning demands, yielding average accuracy improvements of 19.52% and 8.06%, respectively, on the dataset.
@@ 1 Introduction
Vision language models (VLMs) are capable of step-by-step reasoning, presented as Chain-of-Thought (CoT) Wei et al. (2022), for image understanding in question answering (Q&A) across various scientific domains (Chen et al., 2024a; Yue et al., 2024; Wu et al., 2025d), driven by the development of domain-specific benchmarks (Weng et al., 2025; Wang et al., 2025b; Yue et al., 2024; Wang et al., 2025a), step-wise annotations (Wang et al., 2024c; Zhang et al., 2024b; Qi et al., 2024), and domain-tailored VLMs (Yang et al., 2023; Huang et al., 2024; Lan et al., 2025). In contrast, advancements in the financial domain remain limited. This is because, in finance image understanding, the creation of CoT reasoning datasets, which are highly necessary for the evaluation and fine-tuning of VLMs, requires a high level of expertise and domain knowledge and is therefore scarce.
@@
To address these issues, previous literature (Chen et al., 2021) collects Q&A samples from financial documents manually, a process that is prohibitively time-consuming and difficult to scale. As a result, subsequent works propose datasets by merging extensive existing financial datasets (Shah et al., 2022; Xie et al., 2024). These repurposed datasets may introduce accumuated biases and inconsistencies inherited from the original sources. More recently, benchmarks such as (Gan et al., 2025; Kim et al., 2024; Luo et al., 2025; Lu et al., 2025; Li et al., 2025a; Liu et al., 2025d) leverage trained VLMs to generate Q&A pairs augmented with CoTs (Xue et al., 2025). However, these samples may contain factual errors or hallucinations, which can distort VLM evaluation and fine-tuning and thereby undermine their reliability in real-world financial applications, where precision and expert-level understanding are essential.
@@
PyFi-600K offers three key benefits for the interpretability and fine-tuning of VLMs in finance. First, the samples are organized in a pyramid structure, where the required image understanding capability increases progressively — from basic visual perception at the bottom level to complex financial decision-making at the top. This hierarchy, with capability levels ranging from one to six, enables comprehensive evaluation of VLMs across multiple aspects and supports ranking the capabilities required by VLMs in different financial themes. Second, samples across different levels are interconnected through question chains, demonstrating how questions at higher levels can be addressed step-by-step by leveraging information and reasoning from lower-level Q&A samples. Finally, it supports process supervision (Lightman et al., 2023), as each sample is assigned a reward score that reflects the correctness of its reasoning process in solving the financial problem.
@@ 2 Related Work
Financial Benchmark. The evaluation of Vision-Language Models (VLMs) in the finance domain has seen remarkable progress, from early text-based benchmarks (Chen et al., 2021; Shah et al., 2022; Xie et al., 2024) to more recent efforts targeting financial image understanding (Kim et al., 2024; Zhu et al., 2025a; Shu et al., 2025; Li et al., 2025a; Liu et al., 2025d; Rangapur et al., 2025; Zhao et al., 2025; Gan et al., 2025; Luo et al., 2025; Xue et al., 2025; DENG et al., 2025). However, the question-answering datasets of most benchmarks, such as FinMME (Gan et al., 2025), rely on human annotation, making them costly and difficult to scale. Besides, they fail to support comprehensive assessment of VLMs in the financial domain. However, we introduce an automatically synthesized dataset for the fine-grained, hierarchical evaluation of financial visual understanding across diverse image types and financial themes. More related background can be found in Appendix A.
Financial Large Models. Previous efforts primarily centered on text analysis using financial LMs (Huang et al., 2023; Wu et al., 2023; Yang et al., 2023; Xie et al., 2023; Liu et al., 2025c). Recent financial VLMs (Wang et al., 2023; Bhatia et al., 2024; Huang et al., 2024; Park et al., 2025) integrate visual inputs to facilitate step-by-step reasoning, thereby enabling broader financial decision-making. However, these models lack the expertise and interpretability needed for challenging financial themes, a limitation stemming from the absence of detailed, step-wise problem-solving processes grounded in finance-specific knowledge. In contrast, our dataset contains questions that illustrate how to address financial problems through a progression from simple to complex image understanding and from basic to increasingly deep financial knowledge.
Multi-Agent LM Framework. Recent advances in LMs have enabled Agentic AI to excel in real-world applications, with multiple LM-based agents now collaborating to perceive, learn, reason, and act, supporting intricate tasks such as simulation (Lin et al., 2025; ShengbinYue et al., 2025), reasoning (Liu et al., 2024a; Wu et al., 2025a; Du et al., 2024), trading (Li et al., 2025b; Yu et al., 2024; Wu et al., 2025c), and data synthesis (Tang et al., 2024; Liu et al., 2025a; Xuan et al., 2025). However, to our knowledge, no existing method in finance leverages a multi-agent adversarial approach to synthesize data. In contrast, inspired by Generative Adversarial Networks (GANs) Goodfellow et al. (2014), we present the first method that employs adversarial agents to automatically synthesize hierarchical, fine-grained financial datasets.
@@ 3 Framework
Figure 2: Overview of PyFi-adv: a challenger agent competes with a solver agent under the Monte Carlo tree search (MCTS) paradigm to generate question chains that progressively probe deeper capability levels in financial visual reasoning. See the complete chain in Figure 12 in Appendix.
@@ 3.1 Financial Dataset in Pyramid Structure
Therefore, compared to existing manual and automatic evaluation dataset construction methods, we observe the following three advantages of the proposed approach for the finance domain:
@@
Step-wise Reward Score $r$ of each sample reflects its reliability; thus, the scores along a sample chain support process supervision (Lightman et al., 2023) in reasoning with financial image understanding. Using these scores, we enable the training of process reward models that act as verifiers for any financial VLM.
@@ 3.2 Automatic Synthesization with Adversarial Agents
Existing financial datasets rely solely on human annotation, while automated methods from other domains, such as rStar-Math (Guan et al., 2025) for mathematical reasoning, fail to meet the requirements of finance, where interpretability and reliable reasoning are essential, along with increasing complexity, domain expertise, and hierarchical cognition. We achieve this by proposing a multi-agent adversarial mechanism, referred to as PyFi-adv, which constructs question chains, as illustrated in Figure 2.
Specifically, PyFi-adv contains a challenger agent that is supported by an off-and-shelf VLM, denoted as $\psi$, and similarly, a solver agent $\phi$. Inspired by the adversarial mechanism in AlphaGo (Silver et al., 2016), our two agents compete within an MCTS framework to automatically synthesize high-quality reasoning samples. In the finance domain, this approach substantially outperforms existing automatic data augmentation methods discussed in subsection A.3.
@@ 4.1 Setup
Table 1: Evaluation of 15 Pre-trained VLMs on PyFi-600K across 6 capability levels and 5 complexity degrees, conducted on the Evaluation Dataset after Exclusion. This dataset is derived from an initial random selection of 1000 instances from PyFi-600K, followed by an additional filtering step to address potential data leakage. The final test set consists of 301 samples. For simplicity, we only present the results from complexity degrees 3, 4, and 5. The full model names listed in the "Model" column are: GPT-4.1, InternVL3-38B, Claude-opus-4-1-20250805, DeepSeek-VL2, Hunyuan-Large-Vision, ERNIE-4.5-turbo-vl, Moonshot-V1-128k-Vision-Preview, Moonshot-V1-32k-Vision-Preview, Moonshot-V1-8k-Vision-Preview, Qwen3-VL-Plus, Qwen2.5-VL-72B-Instruct, Qwen2.5-VL-32B-Instruct, Qwen2.5-VL-7B-Instruct, and Qwen2.5-VL-3B-Instruct.
Model Overall PP DE CA PR LR DS 33 44 55 GPT-4.1 (OpenAI, 2025) 52.99 86.84 47.64 40.62 57.14 54.84 38.46 52.63 46.55 53.85 InternVL3 (Zhu et al., 2025b) 52.91 76.32 56.37 44.53 43.88 53.23 30.77 41.78 44.83 67.31 Claude-opus (Anthropic, 2024) 64.70 80.92 62.74 66.02 58.16 72.58 32.69 66.12 55.60 73.08 GLM-4.5V (Hong et al., 2025) 74.75 89.47 78.07 65.23 75.51 75.81 46.15 72.37 64.22 73.08 DS-VL2 (Wu et al., 2024b) 45.18 76.32 44.58 36.33 42.35 42.74 19.23 34.54 42.24 46.15 Hunyuan-LV (Sun et al., 2024) 59.72 90.79 58.96 36.72 63.27 69.35 51.92 50.00 57.76 80.77 ER-4.5-TVL (Baidu-ERNIE-Team, 2025) 34.47 34.87 33.96 36.72 25.00 45.16 36.54 38.82 25.43 65.38 MV1-128k (Team et al., 2025b) 54.57 80.92 61.08 39.06 41.84 58.87 38.46 46.71 43.97 57.69 MV1-32k (Team et al., 2025b) 54.40 80.26 61.56 37.11 40.31 62.90 38.46 45.72 43.97 61.54 MV1-8k (Team et al., 2025b) 54.90 80.26 61.79 39.84 42.35 58.06 38.46 48.03 43.97 55.77 Q3-VL-Plus (Yang et al., 2025a) 51.00 77.63 54.25 37.50 41.84 57.26 32.69 45.72 37.50 65.38 Q2.5-VL-72B (Bai et al., 2025) 48.84 70.39 59.91 33.20 35.71 46.77 26.92 36.51 31.03 57.69 Q2.5-VL-32B (Bai et al., 2025) 43.19 60.53 45.28 37.50 36.73 42.74 28.85 39.14 30.60 59.62 Q2.5-VL-7B (Bai et al., 2025) 37.87 54.61 49.53 34.77 18.37 24.19 15.38 31.91 17.67 38.46 Q2.5-VL-3B (Bai et al., 2025) 20.51 36.84 15.57 21.48 16.84 21.77 19.23 23.03 18.97 30.77
Datasets. PyFi-600K dataset consists 603,859 samples and 62,660 sample chains with reward scores of 1. For evaluation, we proportionally sample 1,000 samples from each capability level to construct the test set across all cases. See detail in Appendix B.
Data Leakage Mitigation. To ensure a fair evaluation, we systematically identify and exclude samples that exhibit data leakage. Specifically, any question that can be correctly answered by more than half of the evaluated models without the image is considered leaked. This reduces our test set from 1,000 to 301 samples. Detailed analysis and justification are provided in Appendix D.
@@ 4.3 Main Benefits and Insights
Evaluations in Table 1 and Figure 3 enable a fine-grained analysis of how — and to what extent — VLMs meet the varying capability requirements of financial image understanding across different levels of complexity. The phenomenon of decreasing accuracy as the level increases in Table 1 indicates that existing VLMs can only meet the most basic image understanding in finance and are difficult to be used for complex financial decision-making. For instance, while models like GLM-4.5V perform well on basic perception questions (level 1), their accuracy drops considerably on decision support questions (level 6). This trend is consistent across most models, suggesting that greater efforts are needed to enhance VLMs’ capabilities in the finance domain. Thus, the results in Figure 3 demonstrate the importance of our question-chain fine-tuning: Qwen models learn to make accurate financial decisions by first addressing a series of sub-questions in a progression from simple to complex, a process aligned with human cognition.
@@
Tracing back from Decision Support to Lower-level Q&As, as illustrated in Figure 4, significantly improves the interpretability of VLMs in financial decision-making. In Figure 4, we first collect Level-6 samples that VLMs fail to answer correctly and analyze their question chains by evaluating VLMs on the questions at each lower level in the chain. In the finance domain, addressing a Level-6 financial decision-making question typically requires correctly answering sub-questions from lower levels. Thus, when an error occurs, as indicated by the red cross in the right sub-figure, PyFi-600K enables us to trace the problem-solving chain from DS to LR and eventually to PP, thereby diagnosing how an error in the DE leads to the final mistake. Such a tracing-back approach is crucial for understanding the limitations of VLMs in financial decision-making, as it reveals at which capability level a failure occurs.
@@
Identifying the Most Frequent Failure Capability Levels, as presented in the left part of Figure 4, reveals that the key to improving VLMs’ performance in the financial domain lies in strengthening their capabilities at the CA (Calculation Analysis) level. Specifically, the pyramid-like structure and the question chain of PyFi-600K facilitates detailed error analysis, allowing us to identify specific levels where VLMs struggle. That is, most VLMs, even the advanced Claude-opus-4-1-20250805 and Qwen3-VL-Plus, fails to answer more than 40% of the question from CA, revealing that the calculation analysis is a critical bottleneck for them. More importantly, by comparing the error proportions between six capability levels, we can prioritize which capabilities to enhance first. For instance, since CA has the highest error proportion, focusing on improving calculation analysis skills in VLMs could lead to the most significant overall performance gains in financial decision-making tasks. In addition, this comparison also shows that, in the finance domain, analyzing numerical calculations from images is one of the most critical capabilities for accurate decision-making, even more important than logical reasoning, aligning with conclusions drawn by real-world finance experts.
@@
Addressing Financial Decision-Making with Level-wise CoT, as depicted in Figure 5, is not only reliable but also highly interpretable. Specifically, we compare the responses of Qwen2.5-VL-3B-Instruct and PyFi-QwenVL-3B — the latter fine-tuned on question chains— when addressing the financial problem. The Qwen2.5-VL-3B-Instruct produces a short and disorganized reasoning process, reflecting its limited financial reasoning capability. In contrast, PyFi-QwenVL-3B follows a clear, structured, and logical reasoning flow, which is essential for expert-level performance in finance. For instance, our model first addresses basic PP questions, then proceeds to DE and CA, and finally tackles high-level LR and DS. This level-wise CoT not only improves answer reliability but also provides a transparent and traceable reasoning path that users can easily follow and verify, making it especially valuable in financial contexts where stakes are high. Moreover, experts can inspect this professional decision-making process and provide detailed and direct feedback to further refine the model.
@@
Providing Process Supervision, in the form of a reward score for each sample in our PyFi-600K, enables research into training reliable verifier models, known as process reward models (PRMs), to facilitate checking the correctness of each step generated by financial VLMs. Even though, due to space limitations, this paper does not perform such training, we use the reward scores to filter out low-quality samples during both evaluation and fine-tuning.
Mitigating Hallucinations. As shown in Figure 5 and Figure 4, our support of Level-wise CoT in training effectively mitigates model hallucination by enforcing a grounded reasoning process: high-level conclusions (e.g., DS) are derived strictly from verified low-level findings (e.g., PP and DE), rather than from ungrounded speculation. This structure ensures traceability and verifiability, as every reasoning step can be explicitly traced back to concrete evidence in preceding steps.
5 Conclusion